### v3

# Meromorphic functions whose action on their Julia sets is Non-Ergodic

Abstract.
Nevanlinna functions are meromorphic functions with a finite number of asymptotic values and no critical values. In [KK2] it was proved that if the orbits of all the asymptotic values accumulate on a compact set on which the function acts as a repeller, then the function acts ergodically on its Julia set. In [CJK4], we proved the action of the function on its Julia set is still ergodic if some, but not all of the asymptotic values land on infinity, and the remaining ones land on a compact repeller. In this paper, we complete the characterization of ergodicity for Nevanlinna functions by proving that if all the asymptotic values land on infinity, then the Julia set is the whole sphere and the action of the map there is non-ergodic.
2010 Mathematics Subject Classification:
Primary: 37F10, 30F05; Secondary: 30D05, 37A30
This material is based upon work supported by the National Science Foundation. It is partially supported by gifts from the Simons Foundation (numbers 523341 and 942077) and PSC-CUNY awards and Yuan-Ling Ye’s NSFC Grant No. 12271185. It was also supported by the National Science Foundation under Grant No. 1440140, while the third author was in residence at the Mathematical Sciences Research Institute in Berkeley, California, during the spring semester 2022
1. Introduction
It is a theorem of McMullen [Mc] and Lyubich [Lyu1]) that if $f$ is a rational map of degree greater than one, and if $P(f)$ is its post-singular set (see the formal definition in the next section), one of two things holds: either the Julia set $J(f)$ is equal to the whole Riemann sphere and the action of $f$ is ergodic or, for almost every $z$ in $J(f)$, the spherical distance $d(f^{n}(z),P(f))\rightarrow 0$ as $n\to\infty$; that is, the $\omega$-limit set $\omega(z)$ is a subset of $P(f)$ that varies with $z$.
In the same vein, Bock [Boc] proved a dichotomy theorem for transcendental functions: either the Julia set is the whole sphere and for any set $A$ of positive Lebesgue measure, the orbits of almost all points in $\mathbb{C}$} have infinitely many iterates that land in $A$, or the Julia set is not the whole sphere, and almost every point in the Julia set is attracted to the post-singular set. Unlike the rational case, though, when the Julia set is the whole sphere, it is not known when such functions are ergodic. Misiurewicz [Mis] has shown that when $f(z)=e^{z}$, the Julia set is the whole sphere, but Lyubich [Lyu2], following earlier work [GGS, Dev], has shown that its action is not ergodic. More examples of entire maps that are ergodic can be found in [WZ, CW]. Skorulski,  [S, S1], answered the ergodicity question for meromorphic maps of the form $f(z)=\lambda\tan z$ and $f(z)=\frac{ae^{z^{p}}+be^{-z^{p}}}{ce^{z^{p}}+de^{z^{-p}}}$ that have two asymptotic values.
A more general sort of affirmative result can be found in [KK2]: If the orbits of the singular values either land on or accumulate on a compact repelling set, and if the mapff satisfies an additional growth rate condition, then its Julia set is the whole sphere and it is ergodic.
In this paper, we study the ergodicity question for meromorphic functions whose singular set consists of finitely many asymptotic values and no critical values, the so-called Nevanlinna functions. The dynamical properties of these functions were first investigated in [DK], where it was shown that they share many of the dynamical properties of rational maps. In [CJK4], the authors studied the ergodicity question for a subfamilies of Nevanlinna functions for which some of the asymptotic values have orbits that land on infinity. They proved:
Theorem A.
Let $f$ be a Nevanlinna function. If some, but not all, of the asymptotic values of $f$ have orbits that land on infinity, and if the remaining asymptotic values have $\omega$-limit sets that are compact repellers, then the Julia set of $f$ is the Riemann sphere, and $f$ is ergodic there.
In this paper we complete the characterization of ergodicity for Nevanlinna functions by proving
Main Theorem 1.
Let $f$ be a Nevanlinna function with $N$ asymptotic values, $\lambda_{1},\ldots,\lambda_{N}$. If all $N$ asymptotic values land on infinity, then the Julia set of $f$ is the Riemann sphere, and the action of $f$ is not ergodic there.
In addition, under the same hypotheses for the asymptotic values of $f$, we prove the following theorems.
Main Theorem 2.
For almost every point $z\in\mathbb{C}$, its $\omega$-limit set satisfies:
$\omega(z)=P_{f}=\cup_{i=1}^{N}\{\lambda_{i},f(\lambda_{i}),\ldots,f^{p_{i}-1}(\lambda_{i}),\infty\}.$
Main Theorem 3.
There is no $f$-invariant finite measure absolutely continuous with respect to Lebesgue measure.
Theorem A, together with Main Theorems 1,2 and 3, give a full answer to ergodicity questions for Nevanlinna functions whose Julia set is the whole sphere and whose asymptotic values land either on infinity or on a compact repeller. The questions of whether the conclusions of all three theorems hold for Nevanlinna functions with asymptotic values whose orbits are infinite and whose accumulation sets are unbounded are still open. Another open question is whether there exists a unique $\sigma$-finite $f$-invariant measure absolutely continuous with respect to Lebesgue measure if all the asymptotic values of $f$ have finite orbits.
The paper is organized as follows. The first part of Section 2 contains basic definitions and standard theorems. The remainder of that section is devoted to a detailed review of the properties of Nevanlinna functions. In particular, the asymptotic values divide a neighborhood of infinity into asymptotic tracts on which the function is exponentially contracting. In order to prove the function is non-ergodic, we need to produce disjoint invariant wandering sets in these tracts, each of which has positive Lebesgue measure. This involves estimating the contraction on the asymptotic tracts. In order to make these estimates, we use some classical techniques. We introduce an “auxiliary” variable in each tract which converts it to a half plane in which we can make the estimates. Section 3 contains the heart of the proof of Main Theorem 1. Distinct invariant wandering sets are constructed and their Lebesgue measure is shown to be positive proving the function is not ergodic. The final section contains the proofs of Main Theorems 2 and 3.
2. Preliminaries
2.1. Definitions and standard theorems
A meromorphic function, $f:{\mathbb{C}}\rightarrow\widehat{{\mathbb{C}}}$ is a local homeomorphism everywhere except at the set $S_{f}$ of singular values. In this paper, we will focus on functions whose singular set is finite so that the singular values are isolated. We will assume this throughout. For these functions, the singular values fall into two categories:
Let $v\in\widehat{\mathbb{C}}$ be a singular value and let $V$ be a neighborhood of $v$. Then
  • If, for some component $U$ of $f^{-1}(V)$, there is a $u\in U$ such that $f^{\prime}(u)=0$, then $u$ is a critical point and $v=f(u)\in V$ is the corresponding critical value, or
  • If, for some component $U$ of $f^{-1}(V)$, $f:U\rightarrow V\setminus\{v\}$ is a universal covering map then $v$ is a logarithmic asymptotic value. The component $U$ is called an asymptotic tract for $v$. Any path $\gamma(t)\in U$ such that $\lim_{t\to 1}\gamma(t)=\infty$, $\lim_{t\to 1}f(\gamma(t))=v$ is called an asymptotic path for $v$.
Note that the definition of an asymptotic tract depends on the choice of the neighborhood $V$. If $V_{1},V_{2}$ are punctured neighborhoods of $v$ and $U_{1}$ and $U_{2}$ are unbounded components of their preimages such that $U=U_{1}\cap U_{2}\neq\emptyset$, we call them equivalent asymptotic tracts. For readability, we will not distinguish between “an asymptotic tract” and its equivalence class.
In the proofs of our results we will repeatedly use the Koebe distortion theorems. Many proofs exist in the standard literature on conformal mapping. (See e.g. [Z, Theorem 6.16]), For the reader’s convenience, we state the theorems as we use them without proof.
Theorem 2.1 (Koebe Distortion Theorem).
Let $f:D(z_{0},r)\to\mathbb{C}$ be a univalent function, then for any $\eta<1$,
  1. (1)
    $|f^{\prime}(z_{0})|\frac{\eta r}{(1+\eta)^{2}}\leq|f(z)-f(z_{0})|\leq|f^{\prime}(z_{0})|\frac{\eta r}{(1-\eta)^{2}}$, $z\in D(z_{0},\eta r)$,
  2. (2)
    If $T(\eta)=\frac{(1+\eta)^{4}}{(1-\eta)^{4}}$, $\frac{|f^{\prime}(z)|}{|f^{\prime}(w)|}\leq T(\eta)$, for any $z,w\in D(z_{0},\eta r)$.
  3. (3)
    $|\arg\frac{f^{\prime}(z)}{f^{\prime}(z_{0})}|\leq 2\ln|\frac{1+\eta}{1-\eta}|$, for any $z\in D(z_{0},\eta r)$.
Theorem 2.2.
Let $f:D(z_{0},r)\to\mathbb{C}$ be a univalent function, and $\eta<1$. Then, for any $A,B\subset D(z_{0},\eta r)$,
$\frac{(1-\eta)^{4}}{(1+\eta)^{4}}\frac{m(A)}{m(B)}\leq\frac{m(f(A))}{m(f(B))}\leq\frac{(1+\eta)^{4}}{(1-\eta)^{4}}\frac{m(A)}{m(B)}.$
2.2. Nevanlinna functions
An important tool in studying meromorphic functions with finitely many critical points and finitely many asymptotic values is that they can be characterized by their Schwarzian derivatives.
Definition 1.
If $f(z)$ is a meromorphic function, its Schwarzian derivative is
$S(f)=(\frac{f^{\prime\prime}}{f^{\prime}})^{\prime}-\frac{1}{2}(\frac{f^{\prime\prime}}{f^{\prime}})^{2}.$
The Schwarzian differential operator satisfies the chain rule condition
$S(f\circ g)=S(f)g^{\prime 2}+S(g)$
from which it is easy to deduce that if $f$ is a Möbius transformation, $S(f)=0$, so that $f\circ g$ and $g$ have the same Schwarzian derivative.
In [N], Chap. XI, §3, Nevanlinna, shows how, given a finite set of points in the plane and finite or infinite branching data for these points, this data defines, up to post-composition by a Möbius transformation, a meromorphic function whose topological covering properties are determined by this data. He proves,
Theorem 2.3.
The Schwarzian derivative of a meromorphic function with finitely many critical points and finitely many asymptotic values is a rational function. If there are no critical points, it is a polynomial. Conversely, if a meromorphic function has a rational Schwarzian derivative, it has finitely many critical points and finitely many asymptotic values. If the Schwarzian derivative is a polynomial of degree $m$, then the meromorphic function has $m+2$ asymptotic values and no critical points.
In the literature, meromorphic functions with polynomial Schwarzian are often called Nevanlinna functions (See e.g. [C, EM].)
In this paper, we continue our study of the properties of the dynamical systems these functions generate. Given a Nevanlinna function, we can define the orbit $\{f^{n}{z}\}$ for any point $z\in{\mathbb{C}}$. In general, these orbits are infinite, but if, for some $m\geq 0$, $f^{m}(z)=\infty$, the orbit is finite. Such points are called prepoles of order $m$.
The standard defintions of stable (Fatou) and unstable (Julia) behavior for rational maps carry over to points with infinite orbits; the prepoles are unstable. In [DK], it is proved that the classification of stable behavior for Nevanlinna maps is the same as that for rational maps. In particular, if all the singular points are unstable, the Julia set is the whole sphere.
The results in [KK2] and [CJK4] together show
Theorem.
If $f$ is a Nevanlinna function with $N$ asymptotic values $\lambda_{1},\ldots,\lambda_{N}$, and if for some $0\leq k<N$, $\lambda_{i},i=1,\ldots,k$ are prepoles, [^1] If some $\lambda_{i}=\infty$, it is a ”prepole of order 0”. and if the $\omega$-limit sets of the remaining NkN-k asymptotic values are compact repellers, then the Julia set is the Riemann sphere and $f$ is ergodic there.
In this paper we complete the answer question of the ergodicity of the action of Nevanlinna functions whose Julia set is the whole sphere by proving
Main Theorem 1.
If $f$ is a Nevanlinna function with $N$ asymptotic values $\lambda_{1},\ldots,\lambda_{N}$, and if ALL of them are prepoles, then its Julia set is the Riemann sphere $\widehat{\mathbb{C}}$ and $f$ is not ergodic.
Before getting into the details we sketch the main ideas in the proof (see also [Lyu1, S]). We construct two disjoint sets $A$ and $B$ such that:
  1. (1)
    Both $A$ and $B$ have positive Lebesgue measure;
  2. (2)
    Both sets are wandering and their orbits are disjoint.
Form the sets $X_{A}=\cup_{n\geq 0}f^{-n}(\cup_{n\geq 1}f^{n}(A))$ and $X_{B}=\cup_{n\geq 0}f^{-n}(\cup_{n\geq 1}f^{n}(B))$. They are both invariant and both have positive Lebesgue measure; this proves that the map is non ergodic.
For the construction of $A$ and $B$, we see that the map $f$ has $N$ asymptotic tracts and $N$ asymptotic values each of which is a prepole. Therefore, each asymptotic tract is expansively mapped over all $N$ asymptotic tracts under some finite number of iterations. We fix an asymptotic tract and choose some quadrilaterals that, under an appropriate change of coordinate are squares. We then show that the Lebesgue measure of the set of points whose orbit always lands on these $N$ asymptotic tract is positive and it wanders to $\infty$; these are ”almost” Markov chains because the map is expanding. If we choose the quadrilaterals so that the distance between them is large enough, since the expansion rate on the tracts are different, their orbits will be disjoint.
2.3. The behavior of $f$ in a neighborhood of infinity
The proof of the main theorem depends on a careful study of the behavior of the Nevanlinna function $f$ in a neighborhood of infinity. (See [CJK4, DK, H, L].)
Let $f$ be a Nevanlinna function with Schwarzian derivative $S(f)=2P(z)$, where the degree of the polynomial $P(z)$ is N2N-2 and the leading coefficient of $P(z)$ is the constant $a\in\mathbb{C}^{*}$. The Schwarzian equation is related to the linear equation $w^{\prime\prime}+P(z)w=0$: that is, if $w_{1},w_{2}$ are linearly independent solutions of this equation, then
$S(\frac{aw_{1}+bw_{2}}{cw_{1}+dw_{2}})=2P(z).$
A special case is $N=2$ where the solutions $w_{1},w_{2}$ can chosen as exponentials whose asymptotic tracts are half planes.
This linear equation is often studied by making changes of variable in the asymptotic tracts in order to estimate the behavior of $f$ there. The new variable turns the asymptotic tract into a half plane and the Liouville transformation turns the Schwarzian equation into an approximation of the equation in the case where $N=2$. The interested reader can find more details in the discussions in [CJK4] and [L]. We follow the notation in those discussions.
Denote solutions to the congruence
$\arg a+N\theta\equiv 0\ \ \mod 2\pi$
by $\theta_{i}$, where $1\leq i\leq N$. The critical rays of $f$ are the half lines $L_{i}=te^{i\theta_{i}},t>0$. For a small $\epsilon_{0}>0$, define the sectors
$S_{i}=\{z\ :\ |\arg z-\theta_{i}|\leq\frac{2\pi}{N}-\epsilon_{0}\}$
for $i=1.\ldots,N$. Thus the sector $S_{i}$ contains the critical ray $L_{i}$ and is contained between the critical rays $L_{i-1}$ and $L_{i+1}$. (By convention, all indices are taken modulo $N$). If we denote the boundaries of the sectors by $B_{i}^{\pm}$, depending on the sign of $\arg z-\theta_{i}$, the critical ray $L_{i}$ is contained in a critical wedge $wed_{i}$ of angle $2\epsilon_{0}$ whose boundaries are the rays $B_{i-1}^{+}$ and $B_{i+1}^{-}$.
In each sector $S_{i}$, the function $f$ has a “truncated solution” which grows like $exp(z^{N/2})$. More precisely, let $R>>0$ and $A_{R}=\{z\ :\ |z|>R\}$. Then in $S_{i}$, one can define the auxiliary variable
${\mathcal{Z}}_{i}(z)=\int^{z}_{Re^{i\theta_{i}}}\sqrt{P(s)}ds=\frac{2\sqrt{a}}{N}z^{N/2}(1+o(1)),z\in S_{i}\cap A_{R}.$
We choose the branch of $\sqrt{P(s)}$ so that in the sector $S_{i}$, for any $z$ on the critical ray $L_{i}$, ${\mathcal{Z}}_{i}(z)$ is on the positive real line. The image of $S_{i}$ under ${\mathcal{Z}}_{i}$ contains a sector in the ${\mathcal{Z}}_{i}$-plane,
${\mathcal{S}}_{i}=\{{\mathcal{Z}}={\mathcal{Z}}_{i}(z)\ :\ |\arg{\mathcal{Z}}_{i}|<\pi-\epsilon\}\subset{\mathcal{Z}}_{i}(S_{i})$
for some small $\epsilon>0$ depending on $N$ and $\epsilon_{0}$. In $\mathcal{S}_{i}$, we define
${\mathcal{U}}_{i}=\mathcal{S}_{i}\cap\{{\mathcal{Z}}_{i}\ :\ \Im{\mathcal{Z}}_{i}>c\}\text{ and }{\mathcal{L}}_{i}=\mathcal{S}_{i}\cap\{{\mathcal{Z}}_{i}\ :\ \Im{\mathcal{Z}}_{i}<-c\}$
for some $c>>0$.
Figure 1. The critical lines and sectors for $N=4$
For $z\in S_{i}$ and ${\mathcal{Z}}_{i}={\mathcal{Z}}_{i}(z)$, we use the Liouville transformation to transform the linear equation to one in the variable ${\mathcal{Z}}_{i}$ that has a solution on ${\mathcal{S}}_{i}$ defined by $F({\mathcal{Z}}_{i}(z))=f(z)$. The map $F({\mathcal{Z}}_{i})$ can be approximately expressed as
(1) $f(z)=F({\mathcal{Z}}_{i})=\frac{A_{i}e^{i{\mathcal{Z}}_{i}}+B_{i}e^{-i{\mathcal{Z}}_{i}}}{C_{i}e^{i{\mathcal{Z}}_{i}}+D_{i}e^{-i{\mathcal{Z}}_{i}}}.$
The sets ${\mathcal{U}}_{i}$ and ${\mathcal{L}}_{i}$ are respectively asymptotic tracts for the asymptotic values $B_{i}/D_{i}\text{ and }A_{i}/C_{i}$ of $F({\mathcal{Z}}_{i})$. Therefore, since
$T_{i}={\mathcal{Z}}^{-1}_{i}({\mathcal{U}}_{i})\text{ and }T_{i-1}={\mathcal{Z}}^{-1}_{i}({\mathcal{L}}_{i})$
are mapped by $f$ to punctured neighborhoods of the asymptotic values $\lambda_{i}$ and $\lambda_{i-1}$ of $f$, $\lambda_{i}$ and $\lambda_{i-1}$ are also the asymptotic values of $F$; that is, [^2] Since this holds for each pair of asymptotic values, $\lambda_{i},\lambda_{i-1}$, it is clear both cannot be infinite.
$\frac{B_{i}}{D_{i}}=\lambda_{i}\quad\text{ and }\quad\frac{A_{i}}{C_{i}}=\lambda_{i-1}.$
By equation (1), on the asymptotic tracts ${\mathcal{U}}_{i}$ and ${\mathcal{L}}_{i}$, the map $F({\mathcal{Z}}_{i})$ can be expressed as a composition
[…]
where
$E_{U}({\mathcal{Z}}_{i})=e^{2i{\mathcal{Z}}_{i}},\,\,E_{L}({\mathcal{Z}}_{i})=e^{-2i{\mathcal{Z}}_{i}};$
$M_{i,U}(\xi)=\frac{A_{i}\xi+B_{i}}{C_{i}\xi+D_{i}},\,\text{ and }\,M_{i,L}(\xi)=\frac{A_{i}+B_{i}\xi}{C_{i}+D_{i}\xi}$
Note that $E_{U}$ and $E_{L}$ are infinite to one universal covering maps of ${\mathcal{U}}_{i}$ and ${\mathcal{L}}_{i}$ onto the punctured disk $D^{*}=D^{*}(0,e^{-2c})$. Both the Möbius maps $M_{i,U}$ and $M_{i,L}$ map $D=D^{*}\cup\{0\}$ injectively onto neighborhoods $N_{i}$ and $N_{i-1}$ of the asymptotic values $\lambda_{i}$ and $\lambda_{i-1}$. Thus we obtain factorizations of the truncated solutions in $T_{i}$ and $T_{i-1}$
$f=M_{i,U}\circ E_{U}\circ{\mathcal{Z}}_{i}\mbox{ and }f=M_{i,L}\circ E_{L}\circ{\mathcal{Z}}_{i}.$
By hypothesis, $f^{k_{i}-1}$ and $f^{k_{i-1}-1}$ map $N_{i}$ and $N_{i-1}$ to neighborhoods $P_{i}$ and $P_{i-1}$ of the poles $p_{i}$ and $p_{i-1}$ and so $f$ maps both $P_{i}$ and $P_{i-1}$ to a neighborhood of infinity. Because it is simpler to estimate maps between finite regions, we introduce a transformation that maps a neighborhood of infinity into a neighborhood of the origin. To this end, let $I(z)=1/z$ and set
$f=I\circ g,\text{where }I=\frac{1}{z},\,g(z)=\frac{1}{f(z)}$
so that the map $g(z)$ maps both $P_{i}$ and $P_{i-1}$ into a neighborhood of the origin. The maps
$h_{i,U}(\xi)=g\circ f^{k_{i}-1}\circ M_{i,U}(\xi)\mbox{ and }h_{i,L}(\xi)=g\circ f^{k_{i-1}-1}\circ M_{i,L}(\xi)$
both fix the origin and so map the disk $D(0,e^{-c})$ into a neighborhood of 0.
Figure 2. The decomposition of $h_{i,U}\circ E_{i,U}$ as a map from the auxiliary plane to the dynamic plane
Next, define the following maps on $T_{i}$ and $T_{i-1}$ in the sector $S_{i}$ respectively as:
$\varphi_{i,U}(z)=f^{k_{i}+1}(z)=I\circ h_{i,U}\circ E_{U}({\mathcal{Z}}_{i}),\ \ \varphi_{i,L}(z)=f^{k_{i-1}+1}(z)=I\circ h_{i,L}\circ E_{L}({\mathcal{Z}}_{i}).$
The maps $\varphi_{i,U}$ and $\varphi_{i,L}$ are compositions of a Mobius map, a univalent map and the exponential map, and they map the respective asymptotic tracts $T_{i}$ and $T_{i-1}$ onto a neighborhood $\Omega$ of $\infty$. See figures 2 and 3. Next, for $z\in T_{i}\cup T_{i-1}\subset S_{i}$, we define the maps:
$\Phi_{i}(z)=\begin{cases}\varphi_{i,U}(z),z\in T_{i}\\ \varphi_{i,L}(z),z\in T_{i-1}.\end{cases}$
Figure 3. The map of the asymptotic tract $T_{i}$ (green) and its image under $f^{k_{i}+1}$ (red)
Note that $T_{i}\subset(S_{i}\cap S_{i+1})\cup wed_{i}$. If $z\in S_{i}\cap S_{i+1}$, there are two choices of the auxiliary variables ${\mathcal{Z}}_{i},\ {\mathcal{Z}}_{i+1}$ for $z$. By the choice of the branch of $\sqrt{P(z)}$, ${\mathcal{Z}}_{i}(z)=-{\mathcal{Z}}_{i+1}(z)$ is in the upper half plane. Therefore, if
$\xi=E_{i,U}({\mathcal{Z}}_{i}(z))=E_{i+1,L}({\mathcal{Z}}_{i+1}(z))$
then $h_{i,U}(\xi)=h_{i+1,L}(\xi)$; that is, $\varphi_{i,U}(z)=\phi_{i+1,L}(z)$ so that the map $\Phi_{i}$ is well-defined.
Recall that the fixed constant $c>>0$ is the large constant that defines the sets ${\mathcal{L}}_{i}$ and ${\mathcal{U}}_{i}$. Assume it is chosen large enough so that for all $1\leq i\leq N$,
  1. (1)
    $f^{j}(\lambda_{i})\notin\cup_{i=1}^{N}T_{i},j=0,1\ldots,k_{i}-1$.
  2. (2)
    $h_{i,U}(z),h_{i,L}(z)$ are univalent on the filled disk $D(0,r)=D^{*}(0,r)\cup\{0\}$.
Define the constants
$m_{i}=\frac{1}{|h^{\prime}_{i,U}(0)|},m^{\prime}_{i}=\frac{1}{|h^{\prime}_{i,L}(0)|}\quad 1\leq i\leq N.$
and $m=\min_{i}\{m_{i},m_{i}^{\prime}\},M=\max_{i}\{m_{i},m^{\prime}_{i}\}$.
2.4. Basic calculations
Choose an $\alpha_{0}>c$, and for integers k>0k>0333We use $k$ here as an index and $k_{i}$ as the order of the prepole $\lambda_{i}$. This should not cause confusion., where $N$ is the number of asymptotic values of $f$, define the sequences of real numbers,
$\alpha_{k}=e^{N\alpha_{k-1}}.$
Lemma 2.4.
If $z\in T_{i}$ and ${\mathcal{Z}}_{i}={\mathcal{Z}}_{i}(z)=x+iy$ with $y>\alpha_{0}$, then
(2) $|\Phi_{i}(z)|\geq m_{i}(e^{2y}-2e^{2c}+e^{4c-2y})$
and
(3) $|\Phi_{i}(z)|\leq m_{i}(e^{2y}+2e^{2c}+e^{4c-2y}).$
Proof.
Suppose that $z\in T_{i}$ satisfies ${\mathcal{Z}}_{i}={\mathcal{Z}}_{i}(z)=x+iy$ with $y>c$, then
$\xi=E_{U}({\mathcal{Z}}_{i})=e^{2i{\mathcal{Z}}_{i}}\in D^{*}(0,r),$
$|\xi|=e^{-2y}=\rho r$ and $\rho=e^{2c-2y}\in(0,1)$. Since the map $h_{i,U}$ is univalent on $D(0,r)$ and $h_{i,U}(0)=0$, Koebe’s distortion theorem implies
$\frac{|h^{\prime}_{i,U}(0)|\rho r}{(1+\rho)^{2}}\leq|h_{i,U}(\xi)|\leq\frac{|h^{\prime}_{i,U}(0))|\rho r}{(1-\rho)^{2}}.$
Since $I(z)=1/z$,
$\frac{(1-\rho)^{2}}{\rho r}\frac{1}{|h^{\prime}_{i,U}(0)|}\leq|I(h_{i,U}(\xi))|\leq\frac{(1+\rho)^{2}}{\rho r}\frac{1}{|h^{\prime}_{i,U}(0)|}.$
Note that
$\frac{(1-\rho)^{2}}{\rho r}=\frac{1}{\rho r}-\frac{2}{r}+\frac{\rho}{r}=e^{2y}-2e^{2c}+e^{4c-2y},$
and
$\frac{(1+\rho)^{2}}{\rho r}=e^{2y}+2e^{2c}+e^{4c-2y}.$
Inequalities (2) and  (3) follow, and the proof of the lemma is complete. ∎
Lemma 2.5.
Suppose $z_{1}\in T_{i}$ and $z_{2}\in T_{j}$ where $1\leq i,j\leq N$ (not necessarily distinct) and that $\Phi_{i}(z_{1})\in S_{i^{\prime}}$ and $\Phi_{j}(z_{2})$ are in $S_{j^{\prime}}$. Set ${\mathcal{Z}}_{i}(z_{1})=x_{1}+iy_{1}$ and ${\mathcal{Z}}_{j}(z_{2})=x_{2}+iy_{2}$. Assume that $N_{0}$ is sufficiently large and that for $k>N_{0}$, $y_{1}\geq\alpha_{k}$, $y_{2}\geq y_{1}+2\alpha_{k-1}$. If
(4) $|(\arg(\Phi_{j}(z_{2}))-\theta_{j^{\prime}})|\geq\frac{4}{N\alpha_{k}},\text{ or equivalently }|\sin({\mathcal{Z}}_{j^{\prime}}(\Phi_{j}(z_{2}))|\geq\frac{2}{\alpha_{k}},$
then $\Im({\mathcal{Z}}_{j^{\prime}}(\Phi_{j}(z_{2})))\geq\Im({\mathcal{Z}}_{i^{\prime}}(\Phi_{i}(z_{1})))+2\alpha_{k+1}$.
Proof.
By hypothesis $z_{1}\in T_{i}$ and $z_{2}\in T_{j}$ with $y_{2}>y_{1}>\alpha_{k}>\alpha_{0}$. Inequality (3) implies that
(5) $\Im({\mathcal{Z}}_{i^{\prime}}(\Phi_{i}(z_{1}))\leq|{\mathcal{Z}}_{i^{\prime}}(\Phi_{i}(z_{1}))|\leq\frac{2|a|^{\frac{N}{2}}}{N}(m_{i}(e^{2y_{1}}+2e^{2c}+e^{4c-2y_{1}}))^{\frac{N}{2}}.$
Moreover, by inequalities (2) and (4), when $k$ is large enough,
[…]
Note that $y_{2}>y_{1}+2\alpha_{k-1}$, and $\alpha_{k}=e^{N\alpha_{k-1}}$, so that
(7) $\begin{split}(m_{j}(e^{2y_{2}}-2e^{2c}+e^{4c-2y_{2}}))^{\frac{N}{2}}\frac{2}{\alpha_{k}}&>2(m_{j}(\frac{e^{2y_{1}+4\alpha_{k-1}}-2e^{2c}+e^{4c-2y_{2}})}{e^{2\alpha_{k-1}}})^{\frac{N}{2}}\\ &=2(m_{j}(e^{2y_{1}+2\alpha_{k-1}}-\frac{2e^{2c}-e^{4c-2y_{2}})}{e^{2\alpha_{k-1}}})^{\frac{N}{2}}\\ &>(m_{i}(e^{2y_{1}}+2e^{2c}+e^{4c-2y_{1}}))^{\frac{N}{2}}+N\alpha_{k+1}/|a|^{N/2}\end{split}$
for $k$ large enough, since $\alpha_{k}\to\infty$ and $y_{1}\geq\alpha_{k}$. Therefore
$\Im({\mathcal{Z}}_{j^{\prime}}((\Phi_{j}(z_{2}))))\geq\Im({\mathcal{Z}}_{i^{\prime}}(\Phi_{i}(z_{1})))+2\alpha_{k+1}.$
Define a set of horizontal strips in ${\mathcal{U}}_{i}$, indexed by $k>0$ as
$Hor_{k}^{i}=\{{\mathcal{Z}}_{i}=x+iy\in{\mathcal{S}}_{i}\ :\ \alpha_{k}+2\alpha_{k-1}\leq y\leq\alpha_{k+1}-2\alpha_{k})\}\cap{\mathcal{U}}_{i}.$
The pull-back of these strips
$H_{k}^{i}={\mathcal{Z}}_{i}^{-1}(Hor_{k}^{i})\subset T_{i}.$
are strips in $T_{i}$.
The image $\Phi_{i}(H_{k}^{i})$ is an annular region $A_{i}$ in a neighborhood of infinity of the $z$-plane that, intersects all of the sectors $S_{j}$ for all $j=1,\ldots,N$, and in particular, all the asymptotic tracts $T_{j}$, all the critical lines $L_{j}$ and all the wedges $wed_{j}$. See Figure 3. Thus, the maps ${\mathcal{Z}}_{j}$’s are well-defined for all $1\leq j\leq N$. Define the map $\Psi_{i,j}$ on the ${\mathcal{Z}}_{i}$-plane by the functional equation
${\mathcal{Z}}_{j}\circ\Phi_{i}=\Psi_{i,j}\circ{\mathcal{Z}}_{i}.$
The maps $\Psi_{i,j}$ are infinite to one. To create regions of injectivity, divide each horizontal strip $Hor_{k}^{i}$ into infinitely many rectangles $Rect_{k,n}^{i}$ of width $\pi$ and define $N$ disjoint sub-rectangles $Rect_{j,k,n}^{i}$ in each rectangle $Rect_{k,n}^{i}$ as:
$Rect_{j,k,n}^{i}=\{{\mathcal{Z}}_{i}=x+iy\ :\ x_{j+1}^{i}+n\pi+3/(N\alpha_{k})\leq x\leq x_{j}^{i}+n\pi-3/(N\alpha_{k})\}$
where $x_{j}^{i}$ satisfies $2x_{j}^{i}+\arg(h_{i,U}^{\prime}(0))=-\theta_{j}$.
Remark 2.1.
Under the map ${\mathcal{Z}}_{j}^{-1}\circ\Psi_{i,j}$, each rectangle $Rect_{k,n}^{i}$ maps injectively onto the annulus in the $z$-plane. The vertical lines $x=x_{j}^{i}+n\pi$ in $Hor_{k}^{i}$ are preimages of the critical rays $L_{j}$. The rectangles $Rect_{j,k,n}^{i}$ in each $Rect_{k,n}^{i}$ injectively map onto the disjoint asymptotic tracts in the annulus.
Figure 4. $Hor^{i}_{k}$ for $N=4$
The next lemma states that in the ${\mathcal{Z}}_{i}$-plane, for any point $x+iy$ in these sub-rectangles $Rect_{j,k,n}^{i}$, the inequality (4) in Lemma 2.5 is satisfied. In particular, the sub-rectangles are mapped away from the critical rays and into an asymptotic tract.
Lemma 2.6.
There exists an integer $N_{0}$ such that if $k>N_{0}$, and if ${\mathcal{Z}}_{i}(z)\in Rect_{j,k,n}^{i}$ for some $i$, for each $n$ and for all $j=1,\ldots,N$, then
$|\arg(\Psi_{i,j}(z))|\geq\frac{2}{\alpha_{k}}.$
Proof.
For the proof, fix $i$ and $n$ and, for readability, omit writing them. Set
$M_{j}=\{{\mathcal{Z}}=x_{j}+iy,\,y\geq\alpha_{0}\},$
$L_{j,k}=\Big\{{\mathcal{Z}}=\Big(x_{j+1}+\frac{3}{N\alpha_{k}}\Big)+iy,\,y>\alpha_{0}\Big\}\text{ and }$
$R_{j,k}=\Big\{{\mathcal{Z}}=\Big(x_{j}-\frac{3}{N\alpha_{k}}\Big)+iy,\,y>\alpha_{0}\Big\}.$
Note that the vertical lines $L_{j,k}$ and $R_{j,k}$ contain the respective left and right vertical borders of $Rect_{j,k}$ and the line $M_{j}$ lies in the complementary region between $R_{j,k}$ and $L_{j-1,k}$.
We claim that
[…]
This says that both $\Phi({\mathcal{Z}}^{-1}(L_{j,k}))$ and $\Phi({\mathcal{Z}}^{-1}(R_{j,k}))$ are bounded away from the images of the critical rays $L_{i}$ and $L_{i+1}$; equivalently, the images of $L_{j,k}$ and $R_{j,k}$ under $\Psi$ are bounded away from the real line, as are the points in $Rect_{j,k}$, and proves the lemma.
Figure 5. Rays and curves in lemma 2.6
To prove the claim, first consider the images of the half-lines $R_{j,k}$ and $M_{j}$ under $E_{U}$:
$E_{U}(R_{j,k})=e^{-2y}e^{2i(x_{j}-\frac{3}{N\alpha_{k}}})\mbox{ and }E_{U}(M_{j})=e^{-2y}e^{2i(x_{j})},\ y\geq\alpha_{0}.$
They are line segments meeting at 0 and the angle between them satisfies
$\arg(E_{U}(R_{j,k}))-\arg(E_{U}(M_{j}))=-\frac{6}{N\alpha_{k}}.$
Next, since $h_{U}:D(0,\rho)\to\mathbb{C}$ is conformal, the images $h_{U}(E_{U}(M_{j}))$ and $h_{U}(E_{U}(R_{j,k}))$ are curves with the initial point 0. Let $\gamma_{M}(t)$ and $\gamma_{R}(t)$ be the lines tangent to these curves at 0, respectively. Then
(9) $\begin{split}\gamma_{M}(t)=te^{i(\arg(h_{U}^{\prime}(0))+2x_{j})},\;t\geq 0,\text{ and }\\ \gamma_{R}(t)=te^{i(\arg(h_{U}^{\prime}(0))+2x_{j}-\frac{6}{N\alpha_{k}})},\;t\geq 0.\end{split}$
That is, the angle between $\gamma_{M}(t)$ and $\gamma_{R}(t)$ is equal to $6/(N\alpha_{k})$.
Suppose $\mathcal{Z}$} is on the line $R_{j,k}$, $\xi=E_{i,U}({\mathcal{Z}})$, and $|\xi|=\rho r$, with $|\rho|<e^{2c-2\alpha_{k}}<<1$. Then, if $k$ is large enough, by Koebe’s distortion theorem,
$|\arg(h_{U}(\xi))-(\arg h_{U}^{\prime}(0)+2x_{j}-\frac{6}{N\alpha_{k}}))|<2\ln(\frac{1+\rho}{1-\rho})<\frac{1}{N\alpha_{k}}.$
That is, $h_{U}(\xi)=r_{0}e^{i\theta}$ for some $r_{0}>0$, where
(10) $\theta-(\arg h_{U}^{\prime}(0)+2x_{j})\geq\frac{4}{N\alpha_{k}}.$
If $w=r_{0}e^{i(\arg h_{U}^{\prime}(0)+2x_{j})}$, then
$I(w)=\frac{1}{r_{0}}e^{-i(\arg h_{U}^{\prime}(0)+2x_{j})}\text{ and }\Phi_{i}(z)=I(h_{U}(\xi))=\frac{e^{-i\theta}}{r_{0}}.$
By the definition of $x_{j}$, $\arg(I(w))=\theta_{j}$, and by inequality (10),
$\arg(\Phi_{i}({\mathcal{Z}}_{i}^{-1}(\mathcal{Z})))-\theta_{j}>4/(N\alpha_{k})$. Note that $\mathcal{Z}$} is an arbitrary point in $R_{j,k}$, and by the definition of $\theta_{j}$, $I(\gamma_{M}(t))$, it is mapped to the positive real line under the map $\mathcal{Z}_{j}$. Since $\Psi_{i,j}=\mathcal{Z}_{j}\circ\Phi_{i}$, we conclude
$\arg(\Psi_{i,j}({\mathcal{Z}}))\geq 2/\alpha_{k}.$
One can check, using similar arguments, that at each point $z$ such that ${\mathcal{Z}}(z)$ is on the line $L_{j,k}$,
$\theta_{j+1}-\arg(\Phi_{i}({\mathcal{Z}}_{i}^{-1}({\mathcal{Z}})))\geq 4/(N\alpha_{k}).$
This completes the proof of the claim and hence the lemma. ∎
Lemma 2.7.
Suppose $k>N_{0}$. If ${\mathcal{Z}}_{i}\in Rect^{i}_{j,k,n}$ for some $n$, then $\Psi_{i,j}({\mathcal{Z}}_{i})\in Hor_{k+1}^{j}$.
Proof.
Suppose $k>0$ and ${\mathcal{Z}}_{i}\in Rect^{i}_{j,k,n}$. Then if ${\mathcal{Z}}_{i}=x+iy$,
$\alpha_{k}+2\alpha_{k-1}\leq y\leq\alpha_{k+1}-2\alpha_{k}.$
By inequality (3),
$\Im\Psi_{i,j}({\mathcal{Z}}_{i})<|\Psi_{i,j}({\mathcal{Z}}_{i})|\leq\frac{2\sqrt{|a|}}{N}\Big(m_{i}(e^{2\alpha_{k+1}-2\alpha_{k}}+2e^{2c}+e^{4c+2y})\Big)^{N/2}<\alpha_{k+2}-2\alpha_{k+1}.$
Since ${\mathcal{Z}}_{i}\in Rect^{i}_{j,k,n}$, by lemma 2.6, $|\arg(\Phi_{i,j}({\mathcal{Z}}))|>2/\alpha_{k}$. Because ${\mathcal{Z}}_{i}\in Rect^{i}_{j,k,n}$, $\Im{\mathcal{Z}}_{j}\geq\alpha_{k}+2\alpha_{k-1}$; thus Lemma 2.5 implies
$\Im(\Psi_{i,j}(z)))\geq 2\alpha_{k+1}>\alpha_{k+1}+2\alpha_{k};$
Thus $\Psi_{i,j}(z))\in Hor_{k+1}^{j}$. ∎
3. Disjoint wandering sets of positive Lebesgue measure
To construct the wandering sets it is more convenient to work with the maps $\Psi_{i,j}$ on the auxiliary planes and then pull back. The first step is to estimate the expansion factor.
Proposition 3.1.
Fix the point ${\mathcal{Z}}_{i}^{*}$ in the ${\mathcal{Z}}_{i}$ plane and assume that $\Psi_{i,j}({\mathcal{Z}}_{i}^{*})\in Hor_{k}^{j}$; then
$|(\Psi_{i,j})^{\prime}({\mathcal{Z}}_{i}^{*})|\geq\frac{\alpha_{k}}{4\pi}.$
Proof.
Set ${\mathcal{Z}}_{j}^{*}=\Psi_{i,j}({\mathcal{Z}}_{i}^{*})\in Hor_{k}^{j}$; then $\alpha_{k}+2\alpha_{k-1}\leq\Im({\mathcal{Z}}_{j}^{*})\leq\alpha_{k+1}-2\alpha_{k}$.
Let $D_{k}=D({\mathcal{Z}}_{j}^{*},\alpha_{k}/2)$. By our choices of $c$ and $\alpha_{0}$, the orbits of all the $\lambda_{i}$’s are outside its preimage, ${\mathcal{Z}}_{j}^{-1}(D_{k})$. It follows that there is a univalent branch, $G$ of $\Psi_{i,j}^{-1}$ defined on $D_{k}$. By Koebe’s $\frac{1}{4}$-theorem,
$D\Big({\mathcal{Z}}_{i}^{*},\frac{|G^{\prime}({\mathcal{Z}}_{j}^{*})|}{4}\frac{\alpha_{k}}{2}\Big)\subset G(D_{k}).$
Because $\Psi_{i,j}$ is univalent on $G(D_{k})$, the radius of $G(D_{k})$ is less than $\pi/2$; that is,
$\frac{|G^{\prime}({\mathcal{Z}}_{j}^{*})|}{4}\frac{\alpha_{k}}{2}\leq\frac{\pi}{2}.$
This implies that
$|\Psi_{i,j}^{\prime}({\mathcal{Z}}_{i}^{*})|\geq\frac{\alpha_{k}}{4\pi}.$
Define a family $\mathcal{G}^{i}=\{S^{i}_{m,n};m,n\in\mathbb{Z}\}$, where $S^{i}_{m,n}$ is a square bounded by the straight lines in the ${\mathcal{Z}}_{i}$-plane given by
$x=x^{i}_{N}+n\pi,\;x=x^{i}_{N}+(n+1)\pi,\;y=m\pi,\;y=(m+1)\pi.$
For each $k>0$, let $\mathcal{G}^{i}_{k}\subset\mathcal{G}^{i}$ be the collection of $S_{m,n}^{i}\subset Hor_{k}^{i}$. It will be convenient to use the index $\delta^{i}$ where $S_{\delta^{i}}=S_{m,n}^{i}$. Note that since $x_{N}^{i}$ is fixed, each such square lies between the vertical lines
$V_{n}=V(x^{i}_{N}+n\pi)=\{{\mathcal{Z}}=x^{i}_{N}+n\pi+iy,y>\alpha_{0}\}$
and
$V_{n+1}=V(x^{i}_{N}+(n+1)\pi)=\{{\mathcal{Z}}=x^{i}_{N}+(n+1)\pi+iy),y>\alpha_{0}\}.$
Denote $\mathcal{G}=\cup_{i=1}^{N}\cup_{k>0}{\mathcal{G}}_{k}^{i}$.
For a square $S_{\delta^{i}}\in\mathcal{G}^{i}_{k}$, define the rectangles ${\mathcal{R}}_{\delta^{i}}^{j}=S_{\delta^{i}}\cap Rect_{j,k,n}^{i}$. It follows from the definition of $Hor_{k}^{i}$ that
$\frac{m(S_{\delta^{i}}\setminus(\cup_{j=1}^{N}{\mathcal{R}}_{\delta^{i}}^{j}))}{m(S_{\delta^{i}})}\leq\frac{6}{\alpha_{k}}.$
By lemmas 2.6 and 2.7, for each $j=1,\ldots,N$, $\Psi_{i,j}$ maps the rectangle ${\mathcal{R}}_{\delta^{i}}^{j}$ onto a simply connected domain contained in the horizontal strip $Hor_{k+1}^{j}$.
Let ${\mathcal{U}}^{j}(S_{\delta^{i}})$ denote the union of all the squares $S_{\delta^{j}}$ entirely contained in the interior of $\Psi_{i,j}({\mathcal{R}}_{\delta^{i}}^{j})$:
${\mathcal{U}}^{j}(S_{\delta^{i}})=\cup S_{\delta^{j}}\text{ where }S_{\delta^{j}}\subset{\rm Interior}(\Psi_{i,j}({\mathcal{R}}_{\delta^{i}}^{j}))\subset Hor_{k+1}^{j}.$
. Note that because all the squares have side length $\pi$,
$\rm{dist}(\partial\Psi_{i,j}({\mathcal{R}}_{\delta^{i}}^{j}),\partial{\mathcal{U}}^{j}(S_{\delta^{i}}))\leq\sqrt{2}\pi.$
Otherwise, more squares could be added inside $\Psi_{i,j}(R^{i}_{\delta_{j}})$.
Figure 6. The definition of $\mathcal{P}_{\delta^{i}}^{j}$
For each $j=1,\ldots,N$, let ${\mathcal{P}}_{\delta^{i}}^{j}={\mathcal{P}}^{j}_{\delta^{i}}(S_{\delta^{i}})=\Psi_{i,j}^{-1}({\mathcal{U}}^{j}(S_{\delta_{0}^{j}}))$. Since it is the pullback of a union of squares, the following inequality shows that each ${\mathcal{P}}^{j}_{\delta^{i}}$ is a domain whose area is very close to that of the rectangle ${\mathcal{R}}_{\delta^{i}}^{j}$, see Figure 6. Proposition 3.1 shows that the expansion factor for $\Psi_{i,j}$ on $S_{\delta^{i}}$ is at least $\alpha_{k}/(4\pi)$; this says that
${\rm dist}(\partial{\mathcal{P}}^{j}_{\delta^{i}},\partial{\mathcal{R}}_{\delta^{i}}^{j})<\frac{\rm{dist}(\partial\Psi_{i,j}({\mathcal{R}}_{\delta^{i}}^{j}),\partial{\mathcal{U}}^{j}(S_{\delta^{i}}))}{\frac{\alpha_{k}}{4\pi}}\leq\frac{4\pi^{2}\sqrt{2}}{\alpha_{k}}$
so that the set ${\mathcal{R}}_{\delta^{i}}^{j}\setminus{\mathcal{P}}^{j}_{\delta^{i}}$ is contained in a $4\pi^{2}\sqrt{2}/\alpha_{k}$ neighborhood of $\partial{\mathcal{R}}_{\delta^{i}}^{j}$. Thus for the rectangles ${\mathcal{R}}_{\delta^{i}}^{j}$,
$\frac{m({\mathcal{R}}_{\delta^{i}}^{j}\setminus{\mathcal{P}}^{j}_{\delta^{i}})}{m({\mathcal{R}}_{\delta^{i}}^{j})}\leq\frac{4\sqrt{2}\pi^{2}}{\alpha_{k}}$
and for the full square $S_{\delta^{i}}$,
(11) $\frac{m(S_{\delta_{0}^{i}}\setminus(\cup_{i=j}^{N}{\mathcal{P}}^{j}_{\delta^{i}}))}{m(S_{\delta_{0}^{i}})}\leq\frac{4\sqrt{2}\pi^{2}+6}{\alpha_{k}}.$
Thus each ${\mathcal{P}}^{i}_{\delta_{j}}$ is “almost” equal to the rectangle ${\mathcal{R}}_{\delta^{i}}^{j}$ and $\cup_{i=1}^{N}{\mathcal{P}}^{j}_{\delta^{i}}$ is “almost” equal to the square $S_{\delta^{i}}$. The set ${\mathcal{U}}^{j}(S_{\delta_{0}^{i}})=\Psi_{i,j}(\cup_{j=1}^{N}{\mathcal{P}}^{i}_{\delta_{j}})$ is a union of actual squares, which by Lemma 2.5, lie in $Hor^{j}_{k+1}$.
The process of using the maps $\Psi_{i,j}$ and their inverses to push forward and pull back can be iterated any number of times starting from any square $S_{\delta^{i}}$ in some $Hor_{k}^{i}$. The Lebesgue measure of the resulting pullbacks to $S_{\delta^{i}}$ needs to be estimated. In order to do this, we need to introduce some more notation for iterated maps.
Given $l\in{\mathbb{N}}$, let $\boldsymbol{\iota}=(\delta^{i_{0}},\delta^{i_{1}},\ldots,\delta^{i_{l}})$, and $\boldsymbol{\sigma}=\{j_{0},\ldots,j_{l}\}$ where $i_{0},i_{1},\ldots,i_{l},\\ j_{0},\ldots,j_{l}\in\{1,\ldots,N\}$, and $j_{0}=i_{1},\ldots,j_{l-1}=i_{l}$. Denote the composition $\Psi_{i_{1},i_{2}}\circ\Psi_{i_{0},i_{1}}$ by $\Psi_{i_{0},i_{1},i_{2}}$ and for each $l$, inductively set
$\Psi^{l}=\Psi_{i_{0},i_{1},\ldots i_{l}}.$
Define ${\mathcal{P}}^{\boldsymbol{\sigma}}_{\boldsymbol{\iota}}$ as the set consisting of all points in the square $S_{\delta^{i_{0}}}\subset{\mathcal{Z}}_{i_{0}}$ whose orbit under $\Psi^{l}$ lies in the sequence of quadrilaterals ${\mathcal{P}}_{\delta_{j}}^{i_{j+1}}(S_{\delta^{i}_{j+1}})$, for $j=0,\ldots,l-1$.
Denote the family of indices $({\boldsymbol{\iota}},{\boldsymbol{\sigma}})$ for which ${\mathcal{P}}_{\boldsymbol{\iota}}^{\boldsymbol{\sigma}}$ has a nonempty interior, and is contained in $Hor_{k}^{i_{0}}$, by $\mathcal{I}_{k}^{l}$ . For readability, identify the index $(\boldsymbol{\iota},{\boldsymbol{\sigma}})\in\mathcal{I}_{k}^{l}$ with the set of points in ${\mathcal{P}}_{\boldsymbol{\iota}}^{\boldsymbol{\sigma}}$.
Let $N_{0}$ be chosen as in lemma 2.6 and assume $k>N_{0}$. Set $\mathcal{I}^{l}=\cup_{k\geq N_{0}}\mathcal{I}_{k}^{l}$. By definition, for each ${\mathcal{P}}^{\boldsymbol{\sigma}}_{\boldsymbol{\iota}}\in\mathcal{I}^{l}$, its image under $\Psi^{l+1}$ is a union of squares in $Hor_{k+l+1}^{j_{l}}$.
Set
$Y^{l}=\cup_{(\boldsymbol{\iota},\boldsymbol{\sigma})\in\mathcal{I}^{l}}{\mathcal{P}}_{\boldsymbol{\iota}}^{\boldsymbol{\sigma}}.$
Note that, by definition, $Y^{l+1}\subset Y^{l}$. Let
(12) $Y^{\infty}=\cap_{l=0}^{\infty}Y^{l}.$
Lemma 3.2.
There exists a constant $C>0$ such that if $K=\mathcal{I}^{l}_{k}$, then
$\frac{m(K\setminus Y^{l+1})}{m(K)}\leq\frac{C}{\alpha_{|k|+l+1}}.$
Proof.
By definition, if $K$ is the set defined by $\mathcal{I}^{l}_{k}$, then $\Psi^{l+1}(K)$ is a union of squares in $\cup_{\delta^{j}\in\Delta}S_{\delta^{j}}$ in $Hor_{k+1+l}^{j_{l}}$. Denote this union by $W=\Psi^{l+1}(K)$. Recall that each square $S_{\delta^{j}}$ contains the sets $\mathcal{P}_{\delta^{j}}^{i},i=1,\ldots,N$, each of which is mapped to a square, Therefore,
$\Psi^{l+1}(W\cap Y^{l+1})=\cup_{j_{l}\in\Delta}\cup_{i=1}^{N}{\mathcal{P}}_{S_{\delta^{j_{l}}}}^{i}.$
The union of the orbits of the asymptotic values $\lambda_{i}$ is a finite set. Since the points ${\mathcal{Z}}_{j_{1}}^{-1}(Hor_{k+1+l}^{j_{l}})$ lie in the asymptotic tract of $\lambda_{i}$, we can choose $k$ so that they are far from orbit points. In fact, we can choose it so that there is a neighborhood of ${\mathcal{Z}}_{j_{1}}^{-1}(Hor_{k+1+l}^{j_{l}})$ of radius $2\pi$ disjoint from the orbit points. It follows that the branches of $\Psi^{-(l+1)}$ are well defined on this neighborhood.
Denote the branch that maps $W$ back to $K$ by $\Xi$ and apply the Koebe distortion theorem. It says that for $\zeta,\xi\in W$, there exists a constant $C_{0}$ independent of $\Xi$, such that
$\frac{|\Xi^{\prime}(\zeta)|}{|\Xi^{\prime}(\xi)|}\leq C_{0}.$
Then
$\frac{m(W\setminus Y^{l+1})}{m(W)}=\frac{m(\Xi(\cup(S_{\delta^{j_{l}}}\setminus\cup_{i=1}^{N}{\mathcal{P}}_{\delta_{j_{l}}}^{i})))}{m(\Xi(\cup S_{\delta^{j_{l}}}))}\leq C_{0}^{2}\frac{\sum m(S_{\delta_{j_{l}}}\setminus\cup_{i=1}^{N}{\mathcal{P}}_{\delta^{j_{l}}}^{i})}{\sum m((S_{\delta^{j_{1}}})}\leq\frac{C}{\alpha_{|k|+l+1}},$
for some constant $C=(4\sqrt{2}\pi^{2}+6)C_{0}^{2}>0$. Note that the last inequality follows from Equation(11). ∎
Lemma 3.3.
If $N_{0}>0$ is the integer defined in lemma 2.6, then for any square $S_{\delta^{i}}\in\cup_{k\geq N_{0}}\mathcal{G}^{i}_{k}$, $m(S_{\delta^{i}}\cap Y^{\infty})>0$.
Proof.
Suppose that $S=S_{\delta^{i}}\in Hor^{i}_{k}$ for some $k>N_{0}$. For each $l\geq 0$, set $S^{l}=S\cap Y^{l}$. According to the previous lemma,
$m(S^{l}\setminus S^{l+1})\leq\frac{m(S^{l})C}{\alpha_{|k|+l}}.$
Equivalently,
$m(S^{l+1})\geq m(S^{l})(1-\frac{C}{\alpha_{|k|+l}}).$
Since
$\sum_{l=0}^{\infty}\frac{1}{\alpha_{|k|+l}}\leq\sum_{l=0}^{\infty}\frac{1}{\alpha_{|k|}^{l}}$
and the right side is convergent, it follows that
$0<C_{1}=\prod_{l=0}^{\infty}\Big(1-\frac{C}{\alpha_{|k|+l}}\Big)<\infty.$
Therefore, for $S^{\infty}=S^{0}\cap Y^{\infty}$,
$m(S^{\infty})\geq m(S)\prod_{l=0}^{\infty}\Big(1-\frac{C}{\alpha_{|k|+l}}\Big)\geq C_{1}m(S)>0.$
Now we are ready to complete the proof of Main Theorem 1.
Proof of Main Theorem 1..
Let $k\geq N_{0}$ where $N_{0}$ is the large constant defined in lemma 2.7. Let ${\mathcal{I}}^{l}$ be defined as above. From the definitions of $Hor_{k}^{i_{0}}$, and by lemma 3.3, we can find two squares $S_{\delta^{i_{0}}}\not=S_{\delta^{i_{0}}}^{\prime}\in Hor_{k}^{i_{0}}$ such that
  1. (1)
    $m(S_{\delta^{i_{0}}}\cap Y^{\infty})>0$ and $m(S_{\delta^{i_{0}}}^{\prime}\cap Y^{\infty})>0$;
  2. (2)
    Every pair of points $({\mathcal{Z}},{\mathcal{Z}}^{\prime})\in(S_{\delta^{i_{0}}},S_{\delta^{i_{0}}}^{\prime}),$ satisfies $|\Im{\mathcal{Z}}|-|\Im{\mathcal{Z}}^{\prime}|\geq 2\alpha_{k-1}$.
Let $W_{1}=S_{\delta^{i}}\cap Y^{\infty}$ and $W_{2}=S_{\delta^{i}}^{\prime}\cap Y^{\infty}$. By construction, for any $l\geq 0$,
  1. (1)
    $\Psi^{l}(W_{1})\subset Hor^{i_{1}}_{k+l}$, $\Psi^{l}(W_{2})\subset Hor^{i_{2}}_{k+l}$ for some $i_{1},i_{2}\in\{1,\ldots,N\}$;
  2. (2)
    for all pairs $({\mathcal{Z}}_{i_{l}}\in\Psi^{l}(W_{1}),{\mathcal{Z}}_{i_{2}}\in\Psi^{l}(W_{2}))$, $|\Im{\mathcal{Z}}_{i_{1}}|-|\Im{\mathcal{Z}}_{i_{2}}|\geq 2\alpha_{k+l-1}$.
The first assertion shows that both $W_{1}$ and $W_{2}$ are wandering sets and the second shows that their forward orbits are disjoint. Note that ${\mathcal{Z}}_{i_{0}}$ is analytic, so both ${\mathcal{Z}}_{i_{0}}^{-1}(W_{1}),{\mathcal{Z}}_{i_{0}}^{-1}(W_{2})$ have positive Lebesgue measure.
Let
$E_{1}=\cup_{m=0}^{\infty}f^{-m}(\cup_{n=0}^{\infty}f^{n}({\mathcal{Z}}_{i_{0}}^{-1}(W_{1})))\supset{\mathcal{Z}}_{i_{0}}^{-1}(W_{1})$
and
$E_{2}=\cup_{m=0}^{\infty}f^{-m}(\cup_{n=0}^{\infty}f^{n}({\mathcal{Z}}_{i_{0}}^{-1}(W_{2})))\supset{\mathcal{Z}}_{i_{0}}^{-1}(W_{2}).$
Both have positive Lebesgue measure in the Julia set of $f$. They are mutually disjoint and completely invariant. This implies that $f$ acts non-ergodically on its Julia set completing the proof of the Main Theorem.
4. Main Theorems 2 and 3
Let $P_{f}=\cup_{i=1}^{N}\{\lambda_{i},f(\lambda_{i}),\ldots,f^{p_{i}-1}(\lambda_{i}),\infty\}$ be the post-singular set. As we proved above, $f$ is not ergodic. From the extended dichotomy of Bock [Boc], we know that for almost every point $z\in{\mathbb{C}}$, $\lim_{n\to\infty}d(f^{n}(z),P_{f})\to 0$. This implies that the $\omega$-limit set $\omega(z)\subset P_{f}$ and in terms of the auxiliary variables, it says that for each such $z$, there exists a fixed $i\in\{1,\ldots,N\}$, a sequence $n_{k}\to\infty$, and an $n_{k_{0}}$ such that for all $n_{k}\geq n_{k_{0}}$, $f^{n_{k}}(z)$ is in the asymptotic tract $T_{i}$ and $\Im{\mathcal{Z}}_{i}(f^{n_{k}}(z))\to\infty$.
The set $Y^{\infty}$ was defined in (12) as a subset of $\cup_{i=1}^{N}{\mathcal{Z}}_{i}$ and we showed in lemma 3.3 that it has positive Lebesgue measure. In order to prove Main Theorem 2, which is a result about points in $\mathbb{C}$}, we need to pull $Y^{\infty}$ back to $\mathbb{C}$}. To do this, define $E^{\infty}=\cup_{i=1}^{N}{\mathcal{Z}}_{i}^{-1}(Y^{\infty}\cap{\mathcal{Z}}_{i})$ and set $E=\cup_{n\in{\mathbb{N}}}f^{-n}(E^{\infty})$.
Since $E^{\infty}\subset E$ and $m(Y)>0$, pulling back, $m(E)>0$. In fact, more is true.
Lemma 4.1.
The set $E$ in $\mathbb{C}$} has full Lebesgue measure; that is, $m({\mathbb{C}}\setminus E)=0$.
Proof.
Assume the contrary, $m(\mathbb{C}\setminus E)>0$, and let $z\in{\mathbb{C}}\setminus E$ be a Lebsgue density point. As above, there is an $i\in\{1,\ldots,N\}$ and a sequence $n_{k}$ such that $\Im{\mathcal{Z}}_{i}(f^{n_{k}}(z))\to\infty$ as $n_{k}\to\infty$. Let $k_{0}$ be the smallest integer such that $f^{k_{0}}(z)$ lies in the asymptotic tract $T_{i}$ and $\Im{\mathcal{Z}}_{i}(f^{n_{k_{0}}}(z))>\alpha_{0}$.
Let $l_{k}<\infty$ be a sequence such that
$\alpha_{l_{k}}-2\alpha_{l_{k}-1}\leq\Im{\mathcal{Z}}_{i}(f^{n_{k}}(z))\leq\alpha_{l_{k}+1}-2\alpha_{l_{k}}.$
Note that these inequalities define a set of strips $\widetilde{Hor^{i}_{l_{k}}}$ that is different from, but intersects the strips $Hor^{i}_{n_{k}}$.
Let ${\mathcal{Z}}^{n_{k}}={\mathcal{Z}}_{i}(f^{n_{k}}(z))$, and let $F_{n_{k}}$ be the branch of $f^{-n_{k}}\circ{\mathcal{Z}}_{i}^{-1}$ that maps ${\mathcal{Z}}^{n_{k}}$ to $z$. Let $Q^{n_{k}}$ be the square centered at ${\mathcal{Z}}^{n_{k}}$ with side length $8\alpha_{l_{k}-1}$. [^4] This is not one of the squares in ${\mathcal{G}}^{i}_{n_{k}}$. Then
$m(Q^{n_{k}}\cap(\cup_{k}Hor^{i}_{l_{k}}))\geq\frac{1}{2}m(Q^{n_{k}}).$
Note that $Q^{n_{k}}$ intersects $Hor^{i}_{l_{k}}$ and/or $Hor^{i}_{l_{k}-1}$. The minimum on the left of the above inequality occurs, for example, when ${\mathcal{Z}}^{n_{k}}$ falls on the mid-line of the gap between $Hor^{i}_{l_{k}}$ and $Hor^{i}_{l_{k}-1}$ because the height of the gap is $4\alpha_{l_{k}-1}$.
Let
$\widetilde{Q}^{n_{k}}=\{z\in Q^{n_{k}}\ :\ z\in S\subset Q^{n_{k}}\cap(Hor^{i}_{l_{k}}\cup Hor^{i}_{l_{k}-1})\text{ for squares }S\in\mathcal{G}^{i}\}.$
By the above,
$m(\widetilde{Q}^{n_{k}})\geq\frac{1}{4}m(Q^{n_{k}}).$
Since lemma 3.3 implies that for any square $S\subset\cup_{k}^{\infty}Hor^{i}_{l_{k}}$, there is a constant $C>0$ such that $m(S\cap Y^{\infty})\geq Cm(S)$, it follows that
$m(\widetilde{Q}^{n_{k}}\cap Y)\geq m(\widetilde{Q}^{n_{k}}\cap Y^{\infty})\geq\frac{C}{4}m(Q^{n_{k}}).$
Let $\widetilde{D}_{n_{k}}=D({\mathcal{Z}}^{n_{k}},8\sqrt{2}\alpha_{l_{k-1}})\supset\widetilde{Q}^{n_{k}}$. Recall that $z=F_{n_{k}}({\mathcal{Z}}^{n_{k}})=f^{-n_{k}}\circ({\mathcal{Z}}_{i})^{-1}({\mathcal{Z}}^{n_{k}})$ and set
$A=|(\mathcal{Z}_{i}\circ f^{n_{k_{0}}})^{\prime}(z)|=|(F_{n_{k_{0}}}^{-1})({\mathcal{Z}}^{n_{k}})^{\prime}|>0.$
Let $U_{k}=F_{n_{k}}(\widetilde{Q}^{n_{k}})$ and denote the respective inscribed and circumscribed circles in $U_{k}$ by $D(z,r_{k})$ and $D(z,R_{k})$. Since $F_{n_{k}}$ is univalent on $\widetilde{D}_{n_{k}}$, it has uniform distortion on $\widetilde{Q}^{n_{k}}$ and by Koebe’s theorem there is a constant $B>0$ such that
$\frac{|F_{n_{k}}^{\prime}(\xi)|}{|F_{n_{k}}^{\prime}(\eta)|}\leq B,\quad\forall\;\xi,\eta\in D_{n_{k}},$
and pulling back to the $z$-plane,
$\frac{R_{k}}{r_{k}}\leq B.$
If $\ell=n_{k}-n_{k_{0}}$, then $\Psi^{\ell}={\mathcal{Z}}_{i}\circ f^{\ell}\circ{\mathcal{Z}}_{i}^{-1}$ and proposition 3.1 implies
$|(\Psi^{\ell})^{\prime}(z)|\geq\frac{\alpha_{l_{k}}}{4\pi}.$
Since
$F_{n_{k}}=(\Psi^{\ell}\circ\mathcal{Z}^{i}\circ f^{n_{k_{0}}})^{-1},$
its derivative satisfies
$|F^{\prime}_{n_{k}}(z)|\leq\frac{4\pi}{A\alpha_{l_{k}}}.$
Because the diameter of $\widetilde{Q}^{n_{k}}$ is less than $16\sqrt{2}\alpha_{k-1}$, the diameter of $F_{n_{k}}(\widetilde{Q}^{n_{k}})$ tends to 0, and therefore $R_{k}\to 0$.
Furthermore,
$\frac{m(E\cap D(z,R_{k}))}{m(D(z,R_{k}))}\geq\frac{m(E\cap D(z,R_{k}))}{B^{2}m(D(z,r_{k}))}\geq\frac{m(E\cap U_{k}))}{B^{2}m(U_{k}))}$
$\geq\frac{1}{B^{2}}\frac{m(Y\cap\widetilde{Q}^{n_{k}})}{B^{2}m(\widetilde{Q}^{n_{k}})}\geq\frac{1}{B^{4}}\frac{m(Y\cap\widetilde{Q}^{n_{k}})}{m(Q^{n_{k}})}\geq\frac{C}{4B^{4}}.$
Since this inequality says that
$\lim_{k\to\infty}\frac{m(E\cap D(z,R_{k}))}{m(D(z,R_{k}))}\geq\frac{C}{4B^{4}}>0,$
it implies that the density of $z$ in $E$ is positive, contradicting our assumption that $z$ is a density point of the complementary set ${\mathbb{C}}\setminus E$. Therefore $E$ has full Lebesgue measure. ∎
With this lemma, we can now prove
Main Theorem 2.
For almost every point $z\in\mathbb{C}$,
$\omega(z)=P_{f}=\cup_{i=1}^{N}\{\lambda_{i},f(\lambda_{i}),\ldots,f^{p_{i}-1}(\lambda_{i}),\infty\}.$
Proof.
Since $f$ is not ergodic, for almost every point $z\in{\mathbb{C}}$,
$\lim_{n\to\infty}d(f^{n}(z),P_{f})=0.$
This implies that $\omega(z)\subseteq P_{f}$ for almost every point $z\in{\mathbb{C}}$. To prove the theorem we need to show that $\{z\in{\mathbb{C}},\;\;\omega(z)\subsetneq P_{f}\}$ has zero Lebesgue measure. By lemma 4.1, we only need to show that
$\{z\in E,\;\;\omega(z)\subsetneq P_{f}\}$
has zero Lebesgue measure.
In each plane $\mathcal{Z}_{i}$, let $S=S_{\delta_{i}}$ be a square in $\mathcal{G}_{k}^{i}$ in the strip $Hor^{i}_{k}$. Given $l\in{\mathbb{N}}$, and a $q\in\{1,\ldots,N\}$, let
$K_{l}^{q}=\{{\mathcal{P}}^{\boldsymbol{\sigma}}(S)\ :\ \boldsymbol{\sigma}=\{\delta^{i_{0}},\ldots,\delta^{i_{l}}\},i_{0}\neq m,\ldots,i_{l}\neq m\}.$
That is, $K_{l}^{m}$ consists of all regions $L$ such that none of its successive images under the composition map $\Psi^{l}$ belong to the ${\mathcal{Z}}_{m}$ plane, and whose final image is a square $S^{\prime}$ in a horizontal strip $Hor^{j}_{k+l}$ of some $\mathcal{Z}_{j}$-plane.
Recall that for any $i$, each square $S^{\prime}=Rect^{i}_{k+l}\subset Hor^{i}_{k+l}$ is evenly divided into $N$ rectangles $Rect^{i}_{j,k+l,n}$; this implies that
$m(S^{\prime}\cap Rect^{i}_{j,k+l,n})\leq\frac{1}{N}m(S^{\prime}),$
which in turn implies that
$m(S^{\prime}\cap\cup_{j\neq m}Rect^{i}_{j,k+l,n})\leq\frac{N-1}{N}m(S^{\prime}).$
Fix $q^{\prime}$ and fix $L_{q^{\prime}}={\mathcal{P}}^{\boldsymbol{\sigma}}_{\boldsymbol{\iota}}(S)\in K^{l}_{q^{\prime}}$. The Koebe distortion theorem for the map $\Psi^{l}$ implies that there is a distortion factor $D$ such that for each $q$
$\frac{m(L_{q^{\prime}}\cap(\cup_{L\in K_{l+1}^{q}}L))}{m(L_{q^{\prime}})}\leq D\frac{m(S^{\prime}\cap\cup_{j\neq q}Rect^{j})}{m(S^{\prime})}\leq D\frac{N-1}{N}.$
Because $\Psi^{l}$ is univalent for all $l$, its image is outside a large disk in some ${\mathcal{Z}}^{i}$ plane. Thus, the distortion factor $D$ on each square $S^{\prime}$ of side $\pi$ is close to $1$. Thus we can assume that ND/(N1)<1ND/(N-1)<1; for example
$D=\frac{N+a}{N-a},\text{for some }0<a<\frac{N}{2N-1}$
Therefore, for each $q$
$\frac{m((\cap_{l=0}^{\infty}\cup_{{\mathcal{P}}\in K_{l}^{q}}{\mathcal{P}})\cap S)}{m(S)}=0$
and
$m((\cap_{n=0}^{\infty}\cup_{{\mathcal{P}}\in K_{l}^{q}}{\mathcal{P}})\cap S)=0.$
Finally since $S$ was arbitrary, set
$\mathcal{W}_{i}=\cup_{q=1}^{N}\cup_{k}\cup_{S\in Hor_{k}^{i}}(\cap_{n=0}^{\infty}\cup_{{\mathcal{P}}\in K_{l}^{q}}{\mathcal{P}}).$
Thus $m(\mathcal{W}_{i})=0$, and if $W=\cup_{i=1}^{N}Z_{i}^{-1}(\mathcal{W}_{i})$, then $m(W)=0$; this implies $\cup_{n=1}^{\infty}f^{-n}(W)$ also has zero Lebesgue measure.
To complete the proof, note that $\{z\in E:\omega_{f}(z)\subsetneq P_{f}\}\subset\cup_{n=1}^{\infty}f^{-n}(W)$, so that it has zero Lebesgue measure. ∎
Main Theorem 3.
There is no $f$-invariant finite measure absolutely continuous with respect to Lebesgue measure.
Proof.
Suppose there is an $f$-invariant absolutely continuous measure $\rho$.
Let $\mathcal{W}^{i}(k)=Y^{\infty}\cap\{z\in{\mathcal{Z}}_{i},|\Im z|>\alpha_{k}\}$ and $W_{0}^{i}(k)={\mathcal{Z}}_{i}^{-1}(\mathcal{W}^{i}(k))$. Then
$f^{p_{i}+1}(W_{0}^{i}(k))\subset\cup_{i=1}^{N}W_{0}^{i}(k+1)\subsetneq\cup_{i=1}^{N}W_{0}^{i}(k).$
For each $i=1,\ldots,N$ and $j=1,\ldots,p_{i},$ set $W_{j}^{i}(k)=f^{j}(W_{0}^{i}(k))$. For each $k$, $W_{j}^{i}(k)$ is a bounded set containing $f^{j}(\lambda_{i})$; let $\beta_{k}^{j}$ be the radius of the maximal disk in $W_{j}^{i}(k)$ centered at $f^{j}(\lambda_{i}).$ By lemma 2.7, these disks form a nested sequence whose radii go to zero, so that as $k\to\infty$, the sequences $\beta_{k}^{j}\to 0$.
Let $\mathbf{W}(k)=\cup_{i=1}^{N}\cup_{j=0}^{p_{i}}W_{j}^{i}(k)$ and let $p=\max\{p_{1},\ldots,p_{N}\}$; then, for all pairs $q,n\in\mathbb{N}$,
$f^{(p+1)n+q}(\mathbf{W}(k))\subset\mathbf{W}(k+n).$
Since $E^{\infty}\subset\cup W_{0}^{i}(1)\subset\mathbf{W}(1)$, it follows from lemma 4.1 that
$m({\mathbb{C}}\setminus\cup_{n\in\mathbb{N}}f^{-n}(\mathbf{W}(1)))=0.$
By the absolutely continuity of $\rho$,
$\rho({\mathbb{C}}\setminus\cup_{n\in\mathbb{N}}f^{-n}(\mathbf{W}(1)))=0.$
Moreover, since $\rho$ is also $f$-invariant, there is an $r>0$ such that $\rho(\mathbf{W}(1))=r$. Furthermore, since $f^{k(p+1)}(\mathbf{W}(1))\subset\mathbf{W}(k)$ for each $k$, the invariance implies that $\rho(\mathbf{W}(k))\geq\rho(\mathbf{W}(1))=r$.
Claim: There exists an $r^{\prime}>0$ such that for each $k$,
$\rho(\mathbf{W}(k)\setminus\cup_{i=1}^{N}W_{0}^{i}(k))\geq r^{\prime}.$
If not, for all $r^{\prime}\leq r/2$, there exists a $k$ such that $\rho(\mathbf{W}(k)\setminus\cup_{i=1}^{N}W_{0}^{i}(k))<r^{\prime}$. However, assuming this implies
$\rho(\mathbf{W}(k)\setminus\cup_{i=1}^{N}W_{0}^{i}(k))\geq\rho(\cup_{i=1}^{N}W_{1}^{i}(k))=\rho(f(\cup_{i=1}^{N}W_{0}^{i}(k)))\geq\rho(\cup_{i=1}^{N}W_{0}^{i}(k))\geq r-r^{\prime}\geq r^{\prime},$
which is a contradiction. Thus the claim holds and so for all $k$,
$\rho(\cup_{i}^{N}\cup_{j=1}^{p_{i}}W_{i}^{j}(k))\geq r^{\prime}.$
Furthermore since
$\cap_{k=1}^{\infty}(\cup_{i=1}^{N}\cup_{j=1}^{p_{i}}W_{i}^{j}(k)))=\cup_{i=1}^{N}\{\lambda_{i},\ldots,f^{p_{i}-1}(\lambda_{i})\},$
it follows that
$\rho(\cup_{i=1}^{N}\{\lambda_{i},\ldots,f^{p_{i}-1}(\lambda_{i})\})\geq r^{\prime}$
which contradicts the absolute continuity of $\rho$.